OpenStudy (anonymous):

Given the function g(x) = -4x + 5, find g(-4).

OpenStudy (anonymous):

just replace x in the function by -4 and work it out

hero (hero):

g(x) = -4x+5 g(-4) = -4(-4) + 5 g(-4) = 16 + 5 g(-4) = 21